Consider the following information about a business Dina opened last year: price = $10, quantity sold = 25,000; implicit cost = $155,000; explicit cost = $260,000. What was Dina's economic profit?

1
Expert's answer
2021-11-25T10:18:21-0500

Economic profit =total revenue - explicit cost - Implicit cost


"=250000-260000-155000 \\

=-165000"


"Total \;revenue =Price \times quantity"

"=10\times25000 \\

=250000"